# Multi-scale Graph Autoregressive Modeling: Molecular Property Prediction via Next Token Prediction

### v2

@@ 4.2 Main Results
Unified Takeaway. Taken together, Tables 12 indicate that CamS’s three-stage design (motif mining with single-atom completion $\rightarrow$ causal serialization $\rightarrow$ fine-to-coarse multi-scale concatenation) yields two complementary benefits: (1) general-purpose representation strength, by organizing functional-group semantics, attachment patterns, and topology into an NTP-friendly causal context; and (2) activity-cliff-oriented discriminability, by exposing cliff-driving edits through motif-level abstraction, position-sensitive causal contexts, and redundant multi-scale views. This motivates a targeted interpretability analysis below, where we test whether the model indeed allocates more attention to cliff-differential tokens than to shared scaffolds, and how this preference varies across scale regions.
@@
Task Previous Best Baselines KPGT CamS-LLaMA
Method Score (Ours)
Classification Tasks (AUROC $\uparrow$)
bace GEM 0.857(0.016) 0.855(0.014) 0.870(0.013)
bbbp GEM 0.895(0.024) 0.908(0.012) 0.942(0.015)
clintox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017)
estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050)
metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059)
sider JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016)
toxcast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008)
[…]
AVG GEM 0.825 0.843 0.845
Regression Tasks (RMSE $\downarrow$)
esol GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
freesolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
AVG MolF 1.272 1.175 1.172
Table 2: Performance on MoleculeACE (RMSE). AVG: Average RMSE across all 30 tasks. Ranking: Bold (1st), Underline (2nd), Italic (3rd). Metric: $\downarrow$ denotes lower is better. Column: Previous Best Baselines denotes the best-performing method excluding KPGT; Type: ML (Machine Learning), DL (Deep Learning). Reference & Abbreviations: GMAE: GraphMAE (Hou et al., 2022); MBRT: MoleBERT (Xia et al., 2023); GROV: GROVER (Rong et al., 2020); SVM/RF/GBM: Support Vector Machine / Random Forest / Gradient Boosting Machine (Van Tilborg et al., 2022). Descriptors: E: ECFP (Rogers and Hahn, 2010); M: MACCS (Durant et al., 2002).
Task Previous Best Baselines KPGT CamS-LLaMA
Method Type Score (Ours)
Regression Tasks (RMSE $\downarrow$)
[…]
@@ 4.3 Interpretability: Attention on Activity Cliffs
Interpretability setup and metric. To explain the activity-cliff advantage in Table 2, we test whether CamS-LLaMA preferentially attends to cliff-driving structural differences. Following the MoleculeACE protocol, we construct activity-cliff pairs on the test sets and identify differential vs. shared atom-level activity-cliff point; we then map them to tokens within each scale region of the concatenated CamS sequence (Appendix E). Given the final-layer attention distribution, we compute mean attention on differential tokens (MDTA) and shared tokens (MSTA) within a region and report the relative preference $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100$, evaluated for each scale region and for the full concatenated sequence. We report the full attention metrics (MDTA/MSTA/Rel-DTAP) in Appendix D.
Findings and implications. Table 3 shows three patterns. (1) Overall preference is positive: on the concatenated sequence, Rel-DTAP is consistently >0>0 (14.79% w/o fingerprint; 14.23% w/ fingerprint), indicating that the model, on average, prioritizes cliff-driving differences over shared scaffolds. (2) Preference is scale-dependent: coarse regions (27K/67K) exhibit much stronger preference (24.33%/25.87% w/o fingerprint) than the fewest-merge region (1K, 8.08%), suggesting that coarse motifs provide a more salient interface for highlighting differential substructures, while fine scales contribute weaker but complementary local evidence. (3) Fingerprint acts as a stabilizer rather than a uniform booster: although fingerprint slightly reduces the overall preference, it corrects the anomalous 7K behavior (from -0.18% to 2.10%), consistent with fingerprint providing global guidance that resolves ambiguity at mid-scale tokenizations. Together, these results support the cliff-specific advantage of CamS: multi-scale causal serialization does not merely improve accuracy, but also steers attention toward structurally meaningful differences that drive activity cliffs.
Table 3: Relative Differential-Token Attention Preference (Rel-DTAP). Within each scale region of the concatenated multi-scale sequence, we compute per-molecule $\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%$, then average over all molecules from activity-cliff pairs (anchor and partner counted separately). Positive values mean higher per-token attention on differential tokens than on shared tokens.
Scale region Without FP With FP
1K 8.08% 4.72%
@@ 4.4 Ablation Study
Ablation setup. To isolate the contribution of each core component, we evaluate three variants against the full model (Table 4): (1) w/o FP removes fingerprint injection during fine-tuning; (2) Vocab-1K Only and (3) Vocab-67K Only remove multi-scale concatenation by restricting both pre-training and fine-tuning to a single scale. For fair comparison, single-scale variants keep the pre-training budget matched to the full model (details in Appendix C.1).
Multi-scale concatenation is indispensable. The full model consistently outperforms both single-scale variants across all benchmarks (e.g., MolNet-Cls 0.845 vs. 0.833/0.818; MolNet-Reg 1.172 vs. 1.215/1.329; MolACE 0.624 vs. 0.641/0.649), confirming that multi-scale concatenation is not a minor engineering choice but the key mechanism that lets coarse-scale reasoning build on fine-scale evidence within one causal context. This directly supports our central claim: multi-scale causal serialization is the representation-level ingredient that makes graph structure learnable by a vanilla decoder-only NTP backbone.
Why 1K-only can beat 67K-only under a fixed NTP budget. Under single-scale training, Vocab-67K Only is consistently weaker than Vocab-1K Only, especially on regression (1.329 vs. 1.215 RMSE on MolNet-Reg). We interpret this as an over-compression effect: at 67K, sequences for small molecules become very short and each token packs high compositional complexity, which reduces the number of NTP supervision steps per molecule and makes causal dependency learning harder under limited model capacity and training steps. In contrast, 1K produces longer sequences that provide denser autoregressive supervision and more stable causal credit assignment, albeit at the cost of fragmenting scaffold-level semantics. Concat resolves this trade-off by retaining the dense supervision of fine scales while reintroducing scaffold-level abstraction at coarse scales, without committing to a single “best” vocabulary size.
[…]
Table 4: Ablation study and ranking comparison. Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2. Rank: Global ranking of the specific variant when pooled with the complete set of baseline models. Previous Best: The highest-performing baseline excluding KPGT. Abbreviations: MolF: MolFormer
Method MolNet-Cls ($\uparrow$) MolNet-Reg ($\downarrow$) MolACE ($\downarrow$)
Score Rank RMSE Rank RMSE Rank
CamS-LLaMA (Full) 0.845 1 1.172 1 0.624 1
@@
– Vocab-1K Only 0.833 4 1.215 4 0.641 3
– Vocab-67K Only 0.818 6 1.329 7 0.649 4
@@ 5 Conclusion
We introduced CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction. By unifying motif mining, causal serialization, and fine-to-coarse concatenation, CamS constructs a hierarchical context where global scaffolds are conditioned on dense local evidence. The resulting CamS-LLaMA achieves state-of-the-art performance on MoleculeNet and MoleculeACE, surpassing strong graph baselines while relying on minimal priors. Mechanism analysis confirms that this multi-scale causal design is crucial, as it explicitly drives attention toward cliff-determining structural differences. Our work validates that with the right representation interface, generic autoregressive foundation models can serve as powerful, structure-native engines for molecular science.
@@ Appendix A CamS-tokenizer and Graph-to-Sequence Construction
Overview. This section corresponds to the CamS-tokenizer description in Sec. 3.1 (including Secs. 3.1.1, 3.1.2, and 3.1.3) and the multi-view training-instance construction in Sec. 3.2.1. It supplements the main text with implementation-level details of (1) GraphBPE merge learning and Single-Atom Vocabulary Closure (SAVC)(Algs. 1 and Algs. 2), (2) connection-aware motif vocabulary construction and encoding-time unknown recovery (Algs.3 and Algs.4), (3) deterministic scaffold-rooted BFS serialization (Algs. 5), and (4) fine-to-coarse multi-scale concatenation together with NTP loss masking and length handling (Algs. 6)
A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage (GraphBPE + SAVC)
@@
We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where each node is an atom and each edge is a chemical bond. A GraphBPE merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes (details below). Given a prefix $\mathcal{O}_{\leq k}$, applying merges on $G_{0}$ yields a motif graph $G=(V,E)$, where each motif node $v\in V$ stores the covered atom set $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
@@
Each motif token is stored in a pair form $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$. The first records the fragment SMILES of the core atoms (no connectors), while the second inserts wildcard connector atoms * at all attachment sites together with explicit bond types, making the token identity connection-aware. We materialize these motif pairs by replaying $\mathcal{O}_{\leq k}$ on a tokenizer corpus and inserting connector * atoms for every cut bond (Alg. 3).
@@
A key issue for property prediction is encoding coverage: if a rare but meaningful “single-atom + attachment-pattern” form is absent from the mined vocabulary, it would collapse to [UNK] and erase fine-grained distinctions. SAVC addresses this by (1) enumerating common connection-aware single-atom motifs for typical valence states, and (2) adding an element-wise fallback token X_AltForm to represent rare/atypical forms. At encoding time, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to X_AltForm instead of [UNK] (Alg. 2).
@@
Encoding greedily applies $\mathcal{O}_{\leq k}$ and then maps each resulting motif to its token id. If a motif is still unknown after canonicalization, we recursively split it along its stored merge tree until known motifs are obtained; the recursion stops at single-atom leaves where SAVC backoff applies (Alg. 4).
Algorithm-to-component map.
Alg. 1 learns the ordered merge list $\mathcal{O}$ (Sec. 3.1.1). Alg. 3 materializes the connection-aware motif vocabulary by replaying $\mathcal{O}_{\leq k}$ and inserting connector atoms *. Alg. 2 specifies SAVC enumeration and X_AltForm backoff. Alg. 4 describes encoding-time motif extraction and recursive split for unknown recovery.
A.2 Graph-to-Causal-Sequence Serialization (Traversal and Root Selection)
Motif graph.
After applying $\mathcal{O}_{\leq k}$ to $G_{0}$, we obtain a motif graph $G=(V,E)$ where each node $v\in V$ corresponds to a merged fragment with $\mathrm{atom\_indices}(v)\subseteq V_{0}$. An edge $(u,v)\in E$ indicates that at least one original bond in $G_{0}$ connects atoms across the two fragments.
Scaffold-rooted BFS order.
CamS serializes $G$ into a token sequence by a deterministic scaffold-rooted BFS: we select as root the motif with the largest number of covered atoms (ties broken deterministically, e.g., by smaller node id), and traverse the motif graph breadth-first with a fixed neighbor ordering (e.g., ascending node id). The visitation order defines the final token order, implemented by relabeling motif nodes so that ids $0,1,\dots$ encode the BFS order (Alg. 5). This yields a stable core-to-periphery causal order aligned with prefix-based NTP (Sec. 3.1.2).
Why no explicit edges in the sequence.
We do not append edge/direction tokens: each motif token already includes explicit * attachment sites and bond types, so the attachment pattern is part of the token identity (Alg. 3). The BFS order then provides the autoregressive causal structure at the sequence level.
A.3 Multi-scale Concat Construction and Training-Instance Generation
Prefix slicing and scale definition.
GraphBPE yields a single merge list $\mathcal{O}=(o_{1},\dots,o_{K})$. Each vocabulary scale corresponds to a prefix $\mathcal{O}_{\leq k_{s}}$, which induces a sub-vocabulary $\Sigma_{\leq k_{s}}$ and a tokenizer $\iota_{s}$. This “train once, slice many” construction reuses the same learned merge statistics across all scales (Sec. 3.1.3).
Per-molecule views and fine-to-coarse concatenation.
For a molecule $s$, we generate $S$ single-scale views $\{X_{s_{1}},\dots,X_{s_{S}}\}$ by encoding under tokenizers ordered from fine to coarse. We then build one concatenated view by inserting [CONCAT] as separators and keeping only one global [BOS] and [EOS]. Alg. 6 provides the pseudo-code used in implementation (Sec. 3.2.1).
Loss masking and length handling.
For NTP training, we exclude [BOS], [EOS], and [CONCAT] positions from prediction targets (typically by setting labels to -100 at these indices). Sequences are padded with [PAD] up to the maximum length; if truncation is needed, we keep boundary tokens [BOS] and [EOS] and truncate the interior span consistently across views.
Algorithm 1 GraphBPE merging-operation learning (ordered merge list $\mathcal{O}$)
1:Input: tokenizer corpus $\mathcal{D}$ (RDKit molecules), max iterations $K$, min frequency $f_{\min}$

### v3

@@ 4.2 Results on Downstream Tasks
Unified Takeaway. The results across Tables 1 and 2 reveal a distinct advantage of the CamS framework: it combines general-purpose robustness with fine-grained structural discriminability. The three-stage design—(1) motif mining, (2) causal serialization, and (3) multi-scale concatenation—allows the model to encode both global scaffolds and local variations effectively. By doing so, CamS-LLaMA achieves what KPGT attempts via explicit descriptors: it forces the model to attend to chemically significant substructures, but does so intrinsically through the vocabulary and causal objective rather than extrinsically through handcrafted features. This motivates the targeted interpretability analysis in Section 4.3, where we verify whether the model indeed allocates sharper attention to cliff-differential tokens compared to baselines.
@@
Task Prev. Best KPGT Ours Task Prev. Best KPGT Ours
Method Score Method Score
Classification Tasks (AUROC $\uparrow$) Classification Tasks Cont.
BACE GEM 0.857(0.016) 0.855(0.014) 0.870(0.013) Tox21 Ctxt 0.840(0.028) 0.848(0.017) 0.827(0.028)
BBBP GEM 0.895(0.024) 0.908(0.012) 0.942(0.015) AVG (Cls) GEM 0.825 0.843 0.845
ClinTox GEM 0.905(0.027) 0.946(0.026) 0.935(0.017) Regression Tasks (RMSE $\downarrow$)
Estrogen GEM 0.894(0.048) 0.906(0.034) 0.917(0.050) ESOL GEM 0.803(0.051) 0.804(0.102) 0.761(0.046)
Metstab GROV 0.876(0.046) 0.889(0.057) 0.891(0.059) FreeSolv MolF 2.322(0.613) 2.121(1.025) 2.110(0.959)
SIDER JOAO 0.640(0.012) 0.649(0.011) 0.655(0.016) Lipo GROV 0.625(0.007) 0.600(0.012) 0.645(0.023)
ToxCast GEM 0.733(0.020) 0.745(0.003) 0.724(0.008) AVG (Reg) MolF 1.272 1.175 1.172
Table 2: Performance on MoleculeACE (RMSE $\downarrow$). AVG: Average RMSE across all 30 tasks. Bold: 1st; Underline: 2nd; Italic 3rd. Prev. Best: Best baseline excluding KPGT. Type: ML denotes traditional machine learning; DL denotes deep learning.
Task Prev. Best KPGT Ours Task Prev. Best KPGT Ours
Method Type Score Method Type Score
@@ 4.3 Interpretability: Attention on Activity Cliffs
Interpretability Setup and Metric. To mechanistically explain the activity-cliff superiority observed in Table 2, we investigate whether the CamS-LLaMA attention mechanism inherently prioritizes cliff-driving structural variations. Aligned with the CamS-Tokenizer framework (Section 3.1), we map atom-level differences in activity-cliff pairs to tokens within specific Motif Scales of the concatenated CamS Sequence $\mathbf{X}$. Specifically, for each cliff pair, we identify differential versus shared atoms and project these labels onto the CamS Sequence $\mathbf{X}$. Using the final-layer attention distribution, we compute the Mean Attention on Differential Tokens (MDTA) and Shared Tokens (MSTA) within each scale region $s$. We define the Relative Differential-Token Attention Preference (Rel-DTAP) as:
$\mathrm{Rel\text{-}DTAP}_{s}=\frac{\mathrm{MDTA}_{s}-\mathrm{MSTA}_{s}}{\mathrm{MSTA}_{s}+\epsilon}\times 100\%,$ (7)
averaged over all test pairs. A positive value indicates that the model allocates disproportionately higher attention to the subtle structural edits driving the activity cliff. Full implementation details are in Appendix G.
Findings and Implications. Table 3 reveals three patterns that corroborate our design principles: (1) Intrinsic Structural Discriminability: On the full sequence, Rel-DTAP is consistently positive ($\sim$14%), confirming that even without explicit guidance, the NTP-driven causal objective naturally steers attention toward cliff-driving motifs. (2) Scale-Dependent Hierarchy: Coarse scales (27K/67K) exhibit significantly stronger preference ($\sim$25%) than the fine-grained baseline (1K, $\sim$8%). This validates our hypothesis in Section 3.1 that coarse motifs act as high-level ”semantic anchors,” making structural edits more salient compared to the fragmented 1K view. (3) Fingerprint as a Stabilizer: Fingerprint injection (Section 3.2) does not uniformly boost attention but acts as a corrective stabilizer. It specifically rectifies the anomalous attention pattern at the intermediate 7K scale (from -0.18% to 2.10%), resolving ambiguity where motif granularity may be suboptimal.
Collectively, these findings provide a mechanistic basis for CamS’s dominance on MoleculeACE: the multi-scale architecture explicitly exposes subtle edits as distinct multi-granular tokens, preventing them from being obscured by the local averaging mechanisms common in standard graph encoders.
Table 3: Relative Differential-Token Attention Preference (Rel-DTAP). We report the average Rel-DTAP across all activity-cliff test pairs. A positive value indicates that the model attends more intensively to differential tokens. Scale Region ($s$): Indicates the model’s attention to regions in CamS Sequence corresponding to specific Motif Scales.With FP / Without FP: Indicates whether the fingerprint was injected.
Scale Region ($s$) Without FP With FP
1K (Fine) 8.08% 4.72%
@@ 4.4 Ablation Study
Ablation setup. We evaluate three variants against the full model (summary in Table 4, full breakdown in Appendix H) to isolate core components: (1) w/o FP completely removes fingerprint injection; (2) $1\mathrm{K}$ Only and (3) $67\mathrm{K}$ Only use single fixed scales. Single-scale variants have pre-training and fine-tuning budgets matched to the full model.
Indispensability of Multi-Scale Context.
The full model consistently outperforms single-scale variants (e.g., MolACE 0.624 vs. 0.649), confirming that Cross-Scale Concatenation (Section 3.1) is fundamental. By enabling coarse reasoning to ground on fine evidence, this design validates our central thesis: multi-scale causal serialization is the key representation-level innovation that unlocks vanilla NTP for molecular graphs modeling.
Pitfall of Coarse-Scale Over-Compression.
Single-scale results show that coarse modeling ($s=67\mathrm{K}$) consistently lags behind fine-grained modeling ($s=1\mathrm{K}$). We attribute this to supervision sparsity: over-compressed sequences provide insufficient AR steps for effective learning. The full model circumvents this by concatenation, successfully fusing the dense supervision of fine scales with the coarse-scale scaffold-level abstraction, without committing to a specific “best” resolution.
Fingerprint as the Maximum-Scale Global Token.
The pure-sequence model (w/o FP) proves remarkably robust, particularly in classification where it statistically rivals KPGT (0.838 vs. 0.843). While KPGT resorts to heavy priors to patch MNP’s limitations (e.g., over-smoothing), our use of the fingerprint is fundamentally different: we strategically integrate it as the maximum-scale global token (equivalent to Motif Scale $s=\infty$) to complete the topological hierarchy. Acting as a stabilizer (Section 4.3), it plays a significantly larger role in regression tasks, effectively completing the final link of our multi-scale structural context. Thus, the full model’s superiority stems from a unified causal formulation: $Global+(Fine\to Coarse)$, where the fingerprint serves as the holistic structural anchor driving the fine-to-coarse reasoning.
Table 4: Ablation study. Scores represent the average performance across all tasks in the respective benchmarks from Tables 1 and 2. R: Global ranking in complete baselines.
Method MolNet-Cls MolNet-Reg MolACE
AUROC ($\uparrow$) R RMSE ($\downarrow$) R RMSE ($\downarrow$) R
CamS-LLaMA 0.845 1 1.172 1 0.624 1
@@
– 1K Only 0.833 4 1.215 4 0.641 3
– 67K Only 0.818 6 1.329 7 0.649 4
@@ 5 Conclusion
We presented CamS, a graph-to-sequence interface that enables standard decoder-only Transformers to learn molecular topology via Next Token Prediction. Its FM prototype CamS-LLaMA achieve SOTA on MoleculeNet and MoleculeACE. Our multi-scale causal serialization is fundamental to this success, as mechanistic analysis confirms it explicitly drives attention toward cliff-driving structural differences. Ultimately, CamS validates generic AR FMs as powerful engines for structure-native molecular science given the right representation interface.
@@ Appendix A CamS-Tokenizer and Graph-to-Sequence Construction
Overview. This section supplements the CamS-Tokenizer framework description in Sec. 3.1 and the CamS-LLaMA pipeline in Sec. 3.2. It provides implementation-level details for: (1) Vocabulary Mining: Learning merge operations and constructing the Single-Atom Vocabulary Closure (SAVC) (Algs. 1 and 2); (2) Materialization: Building the connection-aware motif vocabulary (Alg. 3); (3) Per-Scale Encoding: The recursive encoding process with unknown recovery (Alg. 4) and the deterministic Scaffold-Rooted BFS serialization (Alg. 5); (4) Cross-Scale Concatenation: Constructing multi-view training instances for NTP (Alg. 6).
A.1 Tokenizer Vocabulary Mining and Single-Atom Coverage
@@
We start from an RDKit molecule and construct the atom graph $G_{0}=(V_{0},E_{0})$, where nodes are atoms and edges are chemical bonds. A merge list is an ordered sequence of operations $\mathcal{O}=(o_{1},\dots,o_{K})$, where each operation $o_{t}$ is a canonical fragment code extracted from the union of two adjacent nodes. The Motif Scale $s$ corresponds to a prefix length $k_{s}$, such that applying $\mathcal{O}_{\leq k_{s}}$ to $G_{0}$ yields a motif graph $G_{s}=(V_{s},E_{s})$, where each node $v\in V_{s}$ covers a subset of atoms $\mathrm{atom\_indices}(v)\subseteq V_{0}$.
@@
Each motif token is stored in a pair form $(v_{\mathrm{noConn}},v_{\mathrm{withConn}})$. To ensure a deterministic and unique identifier for every substructure, we employ RDKit’s canonicalization routine (Landrum et al., 2021). The first component, $v_{\mathrm{noConn}}$, is the standard canonical SMILES of the fragment. The second component, $v_{\mathrm{withConn}}$, explicitly encodes connectivity information to differentiate chemically identical fragments with different attachment contexts (e.g., a pyridine ring attached at the 2-position vs. the 3-position). Specifically, at each attachment site (severed bond), we insert a dummy atom (wildcard *) preserving the original bond type. To guarantee a unique string representation invariant to atom ordering, we generate the canonical SMILES of this wildcard-augmented fragment. Thus, two motifs are identical if and only if they share the same graph topology, atom types, and attachment configurations. We materialize these pairs by replaying $\mathcal{O}_{\leq k_{s}}$ on the tokenizer corpus (Alg. 3).
@@
A key challenge for predictive encoding is coverage: if a rare “single-atom + attachment-pattern” form is absent from the mined vocabulary, standard tokenizers collapse it to [UNK], creating causal information breakpoints. SAVC addresses this by: (1) enumerating common connection-aware single-atom motifs for all typical valences, and (2) introducing an element-wise fallback token [X_AltForm] to represent rare/atypical forms. During encoding, if a queried motif contains exactly one core atom of element $X$ but the exact form is missing, we back off to [X_AltForm] instead of [UNK] (Alg. 2).
@@
Encoding greedily applies $\mathcal{O}_{\leq k_{s}}$ and maps resulting motifs to token IDs. If a motif is unknown (not in $\Sigma$), we recursively split it along its stored merge tree until known sub-motifs are obtained. The recursion guarantees termination at single-atom leaves, where SAVC back-off logic applies (Alg. 4).
A.2 Graph-to-Causal-Sequence Serialization
Motif Graph Construction.
After applying $\mathcal{O}_{\leq k_{s}}$ to $G_{0}$, we obtain the motif graph $G_{s}=(V_{s},E_{s})$. An edge $(u,v)\in E_{s}$ exists if any original bond in $G_{0}$ connects atoms across fragments $u$ and $v$.
Scaffold-Rooted BFS Order (Intra-Scale Order).
CamS serializes $G_{s}$ into a sequence via a deterministic Scaffold-Rooted BFS. To ensure that the serialization is invariant to the input atom permutation (i.e., canonical), we perform a standard molecule-level canonicalization (via RDKit) before any processing. This establishes a canonical ordering of the original atom indices $0,\dots,N-1$. During BFS on the motif graph $G_{s}$, we define the node_id of a motif $v$ as the minimum canonical atom index among its constituent atoms ($\min_{a\in\mathrm{atoms}(v)}\mathrm{idx}(a)$). We then select the root as the motif with the largest atom count (breaking ties by the smallest node_id) and traverse $G_{s}$ breadth-first. When expanding neighbors, we sort them by their node_id in ascending order. This strategy prioritizes global backbone structure before local substituents, establishing a stable Center-to-Periphery causal order (Alg. 5).
A.3 Multi-Scale Concatenation and Training Views
Scale Definition.
We define a set of Motif Scales $S=\{s_{1},s_{2},\dots,s_{M}\}$ (e.g., $1\mathrm{K},7\mathrm{K},\dots$), each corresponding to a prefix $\mathcal{O}_{\leq s_{i}}$. This “train once, slice many” strategy allows reusing the same learned merge statistics across all scales.
Data Augmentation via Views.
For each molecule, we generate $M$ single-scale views $\{X^{(s_{1})},\dots,X^{(s_{M})}\}$ and one concatenated multi-scale view $\mathbf{X}$. The multi-scale view $\mathbf{X}$ is constructed by concatenating single-scale sequences in Fine-to-Coarse Order (Alg. 6), enabling the model to leverage fine-grained details as context for coarse-grained predictions.
NTP Loss Masking.
For AR training, we compute loss on all motif tokens but exclude special tokens [BOS], [EOS], and [CONCAT] from prediction targets (typically by setting labels to -100).
Algorithm 1 BPE-style Merge Operation Learning (List $\mathcal{O}$)
1:Input: Tokenizer corpus $\mathcal{D}$, max iterations $K$, min frequency $f_{\min}$